OpenStudy (anonymous):

solve (cos theta - cos theta)^2 + (cos theta + cos theta)^2

OpenStudy (anonymous):

@amistre64 could help explain this to me? i think it might be cos^2theta but im not entirely sure

OpenStudy (amistre64):

the word 'solve' is rather vague ... what is it we are wanting to do?

OpenStudy (amistre64):

in general tho (a+b)^2 = a^2 + b^2 +2ab (a-b)^2 = a^2 + b^2 -2ab adding them together just gets us: 2a^2 + 2b^2

OpenStudy (anonymous):

well, its asking us to solve and these are my options sin^2 theta 4cos^2theta 8 cos^2theta

OpenStudy (amistre64):

then what can we conclude?

OpenStudy (amistre64):

since a=b ... (a-a)^2 + (a+a)^2 0 2a what is (2a)^2

OpenStudy (anonymous):

iim not sure, like i did this: (cos theta-cos theta)^2 + (cos theta+cos theta)^2 (0)^2 + (cos theta 2) ^2 0+(cos theta 2 ) ^2 (cos theta 2 )^2 4cos^2 is it the second option?

OpenStudy (amistre64):

4 cos^2 is fine

OpenStudy (anonymous):

really? you sure? i doubt myself lol

OpenStudy (amistre64):

typing out theta is not required , cos(t) works all the same

OpenStudy (anonymous):

ohh kk so my final answer would be my second option , correct? just to be sure

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

thats what we determined yes. restating it over and over again neither strengthens nor diminishes that fact.

OpenStudy (anonymous):

ok thank you very much for helping me out :-)

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks